OpenStudy (anonymous):

The average age of husband and wife was 23 years, 5 years ago at the time of their marriage. At present the average age of the husband, the wife and their child is 20 years. How many years after was the child born? 2, 4, 3, 6, 1

OpenStudy (anonymous):

The child is 4 years old now, so it was born after 1 year. I used excel to find the solution, i'll upload a screenshot if you want....

OpenStudy (anonymous):

thnks bro but i got the ans ..yaa u can upload

OpenStudy (anonymous):

i have a simple answer first of all present average age of the parents is H+W/2=23 then after 5 year the average age of the parents and child is H+5+W+5+C/3 =20 so H+W= 46 and H+W+10+C=60 46+10+C=60 so the answer is C= 60-56=4